OpenStudy (anonymous):

What are the x- and y- intercepts of the following equation: y = 56 – x Answer (0, 56), (56, 0) (-56, 0), (56, 0) (0, 56), (-56, 0) (0, -56), (-56, 0)

OpenStudy (anonymous):

@sauravshakya

OpenStudy (anonymous):

@ktnguyen1 HELP

OpenStudy (anonymous):

when x=0 then y=56-0=56 when y=0 then 0=56-x x=56 So, (0, 56), (56, 0)

OpenStudy (tkhunny):

y = 56 – x Add x x + y = 56 Divide by 56 x/56 + y/56 = 1 Done: (56,0) and (0,56)